Check whether the following system of equations has a solution. (6)

3x+2y+6z+4w =4

x+2y +2z +w =5

x+z+ 3w =3


1
Expert's answer
2020-07-13T18:14:44-0400

Let's consider the matrix of the system:

"\begin{pmatrix}
 3 & 2 & 6 & 4 & |4 \\
 1 & 2 & 2 & 1 & |5 \\
 1 & 0 & 1 & 3 & |3
\end{pmatrix}"

Using Gauss method (= making transformations that do not change the system), we can get this matrix (it still coresponds to the given system):

"\begin{pmatrix}
 2 & 0 & 0 & 9 & |13 \\
 0 & 4 & 0 & -1 & |11 \\
 0 & 0 & -2 & 3 & |7\,\,\, 
\end{pmatrix}"

The rank of the matrix of coefficient is equal to the rank of extended matrix. (An extended matrix is ​​a matrix obtained from a matrix of coefficients to which a column of free terms is attached.) This means that, according to the Rouché–Capelli theorem, the system has a solution.


From this particular matrix we can see that system has the infinite number of solutions. We can leave w as a parameter, and express x, y, z through it. So we get:

"z=-\frac{1}{2}(7-3w) \\
y=\frac{1}{4}(11+w) \\
x=\frac{1}{2}(13-9w)"